OpenStudy (anonymous):

complete the square. x^2+2x=9 explain...

OpenStudy (anonymous):

take half of 2, which is 1, and write \[(x+1)^2=9+1^2=10\]

OpenStudy (anonymous):

now that you have \[(x+1)^2=10\] to solve take the square root and get \[x+1=\pm\sqrt{10}\] so \[x=-1\pm\sqrt{10}\]